OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach

OpenStudy (anonymous):

assuming that he kicks it upward

OpenStudy (anonymous):

vf^2 =vi^2 +2ad 0 = 42^2 + 2(-32.2)d 0= 1764 -64.4d Solve for d